OpenStudy (anonymous):

I'm really horrible when it comes to fractions & variables in the same equation. y/-4 + 5 = 15. can someone show me how to solve for y?

OpenStudy (asnaseer):

you could first multiply both sides by 4 to et rid of the fraction

OpenStudy (asnaseer):

*get

OpenStudy (asnaseer):

\[4(\frac{y}{-4}+5)=4*15\]

OpenStudy (asnaseer):

that will simplify to:\[-y+20=60\]

OpenStudy (anonymous):

-y + 20 = 60 -y = 40 y = -40

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

Its quite easy, you have to get used to fractions in equations. In higher classes, you will have to face much tougher equations with decimals too.

OpenStudy (anonymous):